Your friend claims you can use the change – of – base to write the
expression (In x) as a common logarithm. Is your friend correct?
Explain your reasoning.
1
Expert's answer
2020-06-15T09:58:19-0400

We may use the change of base method (https://en.wikipedia.org/wiki/Logarithm#Change_of_base) and write

"\ln x = \log_{e}x = \dfrac{\log_{10}x}{\log_{10}e} = \log_{10}x\cdot \dfrac{1}{\log_{10}e}" ,

the second multiplier is a constant, so we expressed the natural logarithm as a common logarithm multiplied by a constant.